OpenStudy (anonymous):

9y+4=2+25 i need help...

OpenStudy (anonymous):

what that 2 is suppose to be a 2y

OpenStudy (anonymous):

wait*

Nnesha (nnesha):

can be or cannot

OpenStudy (anonymous):

can

Nnesha (nnesha):

ohh sorry i thought you asking that question to us

OpenStudy (anonymous):

so first you want to get your "y's" to the same side so subtract 2y from both sides and it will cancel out on one side and then 9-2=7 so it would be 7y+4=25 then subtract 4 from each side and it cancels out on one side and then 25-4=21 so then you would have 7y=21, now you divide 7 from both sides and 7 cancels out on one side and the 21 divided by 7= 3 so y=3

OpenStudy (anonymous):

hope it helped.

Nnesha (nnesha):

first you have to subtract 2y both side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i subtract 2y from they 9y as well right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

soo 7y+4=25

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

what i do wrong

Nnesha (nnesha):

now subtract 4 both side

OpenStudy (anonymous):

ooo ok

Nnesha (nnesha):

7y +4 = 25 is right

OpenStudy (anonymous):

sorry i read it wrong, yes 7y+4=25 is correct

OpenStudy (anonymous):

i got 3y=25 ... now do i divide

Nnesha (nnesha):

subtract 4 on right side too

Nnesha (nnesha):

whenever you add or subtract multiply or divide one side also have to do on other side

OpenStudy (anonymous):

3y=21

Nnesha (nnesha):

now divide by 3 both side

OpenStudy (anonymous):

y=21

Nnesha (nnesha):

always divide both side

OpenStudy (anonymous):

ok so what i just did is right?

Nnesha (nnesha):

no divide 21 by 7 also

OpenStudy (anonymous):

o so the answer is y=3

Nnesha (nnesha):

yes

OpenStudy (anonymous):

o ok thank you.